OpenStudy (anonymous):

WILL FAN + MEDAL!!!! What is the exact circumference of a circle with a diameter of 8cm? A. 16 pi cm B. 8 pi cm C. 4 pi cm D. 2 pi cm PLEASE HELP!

OpenStudy (mikurout):

option B

OpenStudy (imstuck):

The circumference is C = pi * d (d for diameter). So you fill in 8 for d and C = 8 pi. See that?

OpenStudy (anonymous):

Circumference of a circle = 2*pi*r Radius is diameter/2 = 8/2 =4 2*pi*4 = ? pi

OpenStudy (anonymous):

thanks @mikurout and @IMStuck

OpenStudy (mikurout):

@AMYCARTER Welcome.